OpenStudy (anonymous):

The heat capacity with Pressure constant is known by the bond Cpm=26.6+0.0241T Calculate the delta H, U, , W, Q. Given that 23 mol of gas and the T initial is 18c, T final 16c, with Pressure constant of 1 atm.

OpenStudy (anonymous):

@Somy?

OpenStudy (anonymous):

@divu.mkr Thank you so much!

OpenStudy (anonymous):

Q= n Cv dT Cp - Cv = R H= n Cp dT W = Pdv P=const V1/T1 = V2/T2 and U = Q+W

OpenStudy (anonymous):

@divu.mkr Should I use the integral\[\int\limits Cpm*dT\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@shrutipande9

OpenStudy (anonymous):

thanks for the effort!

OpenStudy (shrutipande9):

u dont use integral..use the formulas @divu.mkr has told.... to find Q(p)= Cpm x dT

OpenStudy (anonymous):

@shrutipande9 do you have any idea how to solve this? it'll be unambiguously much appreciated!

OpenStudy (anonymous):

@perl ?